arectangle has a perimeter 54".What is the formula for finding the width and the legnth? the width is 1/2 the legnth.
Answer by checkley77(12844)

W=L/2
2L+2W=54
2L+2(L/2)=54
2L+L=54
3L=54
L=54/3
L=18 ANS. FOR THE LENGTH.
W=18/2=9 ANS. FOR THE LENGTH.
PROOF:
2*9+2*18=54
18+36=54
54=54